### v1

@@ Abstract
Keywords:
Latin square orthogonal Latin square net satisfiability solving.
@@ A $k$-net($n$) is a set of $n^{2}$ points and $kn$ lines with the following properties:
$A^{T}A=\begin{bmatrix}nI&J&\cdots&J\\ J&nI&\cdots&J\\ \vdots&\vdots&\ddots&\vdots\\ J&J&\cdots&nI\end{bmatrix}$
@@
Although SAT solvers do not always perform well on mathematical problems, they performed remarkably well in the enumeration of orthogonal pairs of Latin squares of order ten with two relations, and this only required minor modifications to the SAT solver and proof verifier (as described in Section 4). We enumerated all 91 pairs of Latin squares of order ten with two relations in less than 1.75 hours on a desktop computer (see Section 5). Conversely, Delisle’s original backtracking search, written in the programming language C and optimized for speed, used around 488 CPU days in 2010. While some of the improvements in speed of our SAT-based search is undoubtedly from the improvements in computer hardware, this alone cannot account for an over $6000\times$ speedup in CPU time, demonstrating the computational efficiency of SAT solvers on problems of this type.
@@ 2 Background
For example, Proposition 1 implies that every nontrivial relation in a 4-net(10) is complementable to one of type $[4,4,4,4]$, because every nontrivial relation in a 4-net(10) is complementable to a nonempty relation with at most 5 lines in three classes. Now consider the case of a 4-net(10) with two linearly independent nontrivial relations. Note that the sum of two nontrivial relations will produce a third relation. Although this relation will not be linearly independent, it will be nontrivial, and therefore by Proposition 1 complementable to a relation of the type $[4,4,4,4]$. The following proposition constrains the manner in which two linearly independent relations $R_{1}$ and $R_{2}$ will intersect.
@@
$[[1,3,3],[1,3,3],[1,3,3],[1,3,3]],$
$[[1,3,3],[1,3,3],[1,3,3],[2,2,2]],$
$[[1,3,3],[1,3,3],[2,2,2],[2,2,2]],$
$[[1,3,3],[2,2,2],[2,2,2],[2,2,2]],$
$[[2,2,2],[2,2,2],[2,2,2],[2,2,2]].$
Furthermore, Delisle used a counting argument to rule out cases 2 and 4 [10, pg. 17]. Thus, Delisle’s search focused on cases 1, 3, and 5. Interestingly, using our approach a SAT solver rules out cases 2, 3, and 4 in a few seconds each. Given case 3 was ruled out by Delisle using 23 days of compute time, the SAT solver shows its strength at uncovering contradictions by ruling out case 3 relatively quickly. At the end of Section 5 we present a less computational argument ruling out case 3 that was found independently from the SAT computation.
@@ 2.1 Delisle’s symmetry breaking
    Permutation of rows of $A$ and $B$ preserving the row equivalence classes.
@@
    Permutation of columns of $A$ and $B$ preserving the column equivalence classes.
@@
    Permutation of the symbols of $A$ preserving the $A$-symbol equivalence classes.
@@
    Permutation of the symbols of $B$ preserving the $B$-symbol equivalence classes.
@@
Suppose $(A,B)$ is a minimal pair with the symbols in the first column of $A$ not in sorted order within the rows of each row equivalence class. Using row permutations, permute the rows of $(A,B)$ to form $(A^{\prime},B^{\prime})$ such that the rows of the first column of $A^{\prime}$ are now sorted within the rows of each row equivalence class. Consider the first entry $A^{\prime}_{i,0}$ of $A^{\prime}$ that has changed after applying these permutations. (This will also be the first entry of $B^{\prime}$ that has changed, since the same permutations are applied to $A$ and $B$.) Because $A^{\prime}_{i,0}<A_{i,0}$ and $(A_{j,0}^{\prime},B^{\prime}_{j,0})=(A_{j,0},B_{j,0})$ for all $j<i$ we have $(A^{\prime},B^{\prime})<(A,B)$ in contradiction to $(A,B)$ being minimal. ∎
@@
By Proposition 7 if $(A,B)$ is a minimal pair then $A_{0,0}$ must be in $\{0,1,4,7\}$ in case 1 and in $\{0,2,4,6\}$ in cases 3–5. Similarly, by Proposition 8 if $(A,B)$ is a minimal pair then $B_{0,0}$ must be in $\{0,1,4,7\}$ in case 1 and in $\{0,2,4,6\}$ in cases 3–5. Delisle [10, pg. 18] gives the possibilities for the values of $(A_{0,0},B_{0,0})$ in case 1, and the only ones which are in $\{0,1,4,7\}\times\{0,1,4,7\}$ are $(0,0)$, $(1,1)$, $(4,4)$, and $(7,7)$. Similarly, [10, pg. 20] gives the possibilities for the values of $(A_{0,0},B_{0,0})$ in case 3 (and these are identical in cases 4 and 5), and the only ones which are in $\{0,2,4,6\}\times\{0,2,4,6\}$ are $(0,0)$, $(2,2)$, $(4,4)$, and $(6,6)$. ∎
@@ 3.1 Latin square encoding
$x_{i}\mathbin{\rightarrow}s_{i}\quad\text{and}\quad s_{i-1}\mathbin{\rightarrow}s_{i}$
@@ 3.2 Orthogonality encoding
$(A_{ij}=k\land A_{i^{\prime}j^{\prime}}=k\land B_{ij}=l\land B_{i^{\prime}j^{\prime}}=l)\mathbin{\rightarrow}(i=i^{\prime})$
for $0\leq i,j,i^{\prime},j^{\prime},k,l<n$ (cf. Zhang [22, Lemma 1]). Taking the contrapositive and writing this using the variables $A_{ijk}$ and $B_{ijk}$, this is equivalent to the clauses $\lnot A_{ijk}\lor\lnot A_{i^{\prime}j^{\prime}k}\lor\lnot B_{ijl}\lor\lnot B_{i^{\prime}j^{\prime}l}$ where $0\leq i,j,i^{\prime},j^{\prime},k,l<n$ with $i\neq i^{\prime}$. This encoding of orthogonality uses $O(n^{6})$ clauses of length 4. An alternative encoding of orthogonality that performs better in practice [22, Lemma 2] uses $O(n^{4})$ clauses of length 3 and $n^{3}$ new auxiliary variables. To describe this orthogonality encoding, we follow the derivation of Bright, Keita, and Stevens [7] based on a composition square.
@@
$(A_{ijk}\land B_{ijl})\mathbin{\rightarrow}Z_{ilk}$
for $0\leq i,j,k,l<n$. In fact, from $A=ZB$ we also derive the similar clause $(Z_{ilk}\land B_{ijl})\mathbin{\rightarrow}A_{ijk}$, and from $B=Z^{-1}A$ we derive $(Z_{ilk}\land A_{ijk})\mathbin{\rightarrow}B_{ijl}$. These last two types of clauses are technically logically redundant, but in practice they improve the performance of the SAT solver as they allow the solver to make additional useful propagations.
@@ 3.3 Relation encoding
Similarly, Delisle defines an equivalence class on symbol pairs $(s,t)$ where $s$ is a symbol of the first Latin square $A$ and $t$ is a symbol of the second Latin square $B$ using a labelling function $\operatorname{ST\_CLASS}$. Representing $s$ and $t$ as integers in $\{0,\dotsc,9\}$, note that line $s+20$ of the net corresponds to symbol $s$ in the first Latin square, and line $t+30$ of the net corresponds to symbol $t$ in the second Latin square. Concretely, $\operatorname{ST\_CLASS}(s,t)$ is defined to be
@@
$\operatorname{RC\_CLASS}(i,j)=\operatorname{ST\_CLASS}(A_{ij},B_{ij})\text{ for all $0\leq i,j\leq 9$}.$
That is, if $(A_{ij},B_{ij})=(s,t)$ then $\operatorname{RC\_CLASS}(i,j)=\operatorname{ST\_CLASS}(s,t)$. We encode this condition directly into our SAT encoding. As we encode each case 1–5 separately, we can assume the forms of $R_{1}$ and $R_{2}$ are known, and therefore the values of $\operatorname{RC\_CLASS}(i,j)$ and $\operatorname{ST\_CLASS}(s,t)$ are known in advance for all $0\leq i,j,s,t\leq 9$. We encode the relation constraint in contrapositive form: for all $i$, $j$, $s$, $t$ with $\operatorname{ST\_CLASS}(s,t)\neq\operatorname{RC\_CLASS}(i,j)$ we enforce that $(A_{ij},B_{ij})\neq(s,t)$, i.e., $A_{ij}\neq s$ or $B_{ij}\neq t$. In Boolean logic, this becomes the clauses
$\lnot A_{ijs}\lor\lnot B_{ijt}\text{ for all $0\leq i,j,s,t\leq 9$ with $\operatorname{RC\_CLASS}(i,j)\neq\operatorname{ST\_CLASS}(s,t)$}.$
@@ 3.4 Symmetry breaking
$\bigwedge_{\begin{subarray}{c}0\leq k,l\leq 9\\ k>l\end{subarray}}(A_{1,0,k}\mathbin{\rightarrow}\lnot A_{0,1,l}).$
@@
$\bigwedge_{\begin{subarray}{c}0\leq k,l,m\leq 9\\ k\geq l\end{subarray}}\bigl{(}(A_{1,0,m}\land A_{0,1,m}\land B_{1,0,k})\mathbin{\rightarrow}\lnot B_{0,1,l}\bigr{)}.$
@@
$\bigwedge_{\begin{subarray}{c}i\in R^{\prime}\\ 0\leq l<k\leq 9\end{subarray}}(A_{i,0,k}\mathbin{\rightarrow}\lnot A_{i+1,0,l}).$
@@
$\bigwedge_{\begin{subarray}{c}j\in C^{\prime}\\ 0\leq l<k\leq 9\end{subarray}}(A_{0,j,k}\mathbin{\rightarrow}\lnot A_{0,j+1,l}).$
@@
$\bigwedge_{\begin{subarray}{c}s\in S^{\prime}\\ 0\leq i^{\prime}<i\leq 9\end{subarray}}(A_{i,0,s}\mathbin{\rightarrow}\lnot A_{i^{\prime},0,s+1}).$
@@
$\bigwedge_{\begin{subarray}{c}t\in T^{\prime}\\ 0\leq i^{\prime}<i\leq 9\end{subarray}}(B_{i,0,t}\mathbin{\rightarrow}\lnot B_{i^{\prime},0,t+1}).$
@@
$\bigwedge_{\begin{subarray}{c}0\leq k,l\leq 9\\ k>l\end{subarray}}(A_{1,0,k}\mathbin{\rightarrow}\lnot B_{1,0,l}).$
@@
$\bigwedge_{\begin{subarray}{c}0\leq k,l,m\leq 9\\ k>l\end{subarray}}\bigl{(}(A_{1,0,m}\land B_{1,0,m}\land A_{2,0,k})\mathbin{\rightarrow}\lnot B_{2,0,l}\bigr{)}.$
@@ 4.1 Exhaustive enumeration
$\lnot\Bigl{(}\bigwedge_{0\leq i,j\leq 9}(A_{i,j,S_{i,j}}\land B_{i,j,T_{i,j}})\Bigr{)},$
@@ 5 Results
We now discuss our computational results enumerating all 4-nets(10) with at least two nontrivial relations. Our results were run on an Intel i7 CPU running at 2.8 GHz and using the SAT solver CaDiCaL 1.9.4 using up to 250 MiB of memory. Our scripts are freely available at https://github.com/curtisbright/Delisle-MOLS.
We use a Python script to generate SAT instances in each of the five possible forms (cases 1–5) following the encoding described in Section 3. In order to mitigate the effect of randomness in the search, each case was independently solved 45 times, each time with a different random seed. The differing random seeds ensure that each instance of CaDiCaL will make different choices during the solving process. A tabular summary of the results of these trials is provided in Table 1, and a box plot of the running times is provided in Figure 1.
The SAT solver determined that cases 2, 3, and 4 all had no solutions and these cases were always solvable in a few seconds. It is interesting to note that Delisle [10] ruled out cases 2 and 4 theoretically using a counting argument, but despite the fact we did not explicitly use this fact in the SAT encoding, the SAT solver quickly proved unsatisfiability on its own. The fact that the solver also quickly ruled out case 3 suggested to us that case 3 was also resolvable using a counting argument, and we were successful in finding one (included at the end of this section). Unfortunately, the proofs produced by the SAT solver, while logically correct, are not intended to be human-readable and the contents of the proofs did not provide us with any mathematical insight.
@@
Figure 1: A box plot visualization of the running times of the 45 instances solved in each case. The left plot shows the results for cases 1 and 5. The right plot shows the results for cases 2, 3, and 4.
@@
The fact that the SAT solver was able to quickly rule out case 3 inspired us to look for a counting argument that could rule out this case. We were successful using an approach similar to the arguments used by Gill and Wanless to count the number of points of certain type in a net [12, Thm. 3.1].
@@
The counting argument in the proof of Theorem 5.1 provides some theoretical conditions that were speculated on by Delisle in their original work:

### v2

@@ A $k$-net($n$) is a set of $n^{2}$ points and $kn$ lines with the following properties:
$A^{T}A=\begin{bmatrix}nI&J&\cdots&J\\ J&nI&\cdots&J\\ \vdots&\vdots&\ddots&\vdots\\ J&J&\cdots&nI\end{bmatrix}$ (1)
@@
SAT solvers do not always perform well on mathematical problems, particularly when there is underlying structure in the problem unknown to the solver. As an example, integer factorization can be reduced to the SAT problem, but SAT solvers cannot factorize integers nearly as efficiently as algorithms specifically designed to do so using number theoretic concepts. Augmenting SAT solvers with number theoretic reasoning improves their performance, but even still they only outperform traditional algorithms when extra information is available such as random bits of the prime factors [1]. Despite SAT solvers having no efficiency guarantee, they did perform remarkably well in the enumeration of orthogonal pairs of Latin squares of order ten with two relations. Only minor modifications to the SAT solver and proof verifier were required, as described in Section 4. We enumerated all 91 pairs of Latin squares of order ten with two relations in less than 1.75 hours on a desktop computer (see Section 5). Conversely, Delisle’s original backtracking search, written in the programming language C and optimized for speed, used around 488 CPU days in 2010. While some of the improvements in speed of our SAT-based search is undoubtedly from the improvements in computer hardware, this alone cannot account for an over $6000\times$ speedup in CPU time, demonstrating the computational efficiency of SAT solvers on problems of this type.
@@ 2 Background
For example, Proposition 1 implies that every nontrivial relation in a 4-net(10) is complementable to one of type $[4,4,4,4]$, because every nontrivial relation in a 4-net(10) is complementable to a nonempty relation with at most 5 lines in three classes. Now consider the case of a 4-net(10) with two linearly independent nontrivial relations. Note the sum of two nontrivial relations $R_{1}$ and $R_{2}$ will be a third relation, and because the relations are over $\mathbb{F}_{2}$, the sum is the symmetric difference $(R_{1}\cup R_{2})\setminus(R_{1}\cap R_{2})$. Although this third relation will not be linearly independent to $R_{1}$ and $R_{2}$, it will be nontrivial, and therefore by Proposition 1 complementable to a relation of the type $[4,4,4,4]$. The following proposition constrains the manner in which two linearly independent relations intersect.
@@
$[[1,3,3],[1,3,3],[1,3,3],[1,3,3]],$ (case 1)
$[[1,3,3],[1,3,3],[1,3,3],[2,2,2]],$ (case 2)
$[[1,3,3],[1,3,3],[2,2,2],[2,2,2]],$ (case 3)
$[[1,3,3],[2,2,2],[2,2,2],[2,2,2]],$ (case 4)
$[[2,2,2],[2,2,2],[2,2,2],[2,2,2]].$ (case 5)
Furthermore, Delisle used a counting argument to rule out cases 2 and 4 [11, pg. 17]. Thus, Delisle’s search focused on cases 1, 3, and 5. Interestingly, using our approach a SAT solver rules out cases 2, 3, and 4 in a few seconds each. Given case 3 was ruled out by Delisle using 23 days of compute time, the SAT solver shows its strength at uncovering contradictions by ruling out case 3 quite quickly. At the end of Section 5, we also give an argument using no search and a short computation that rules out case 3. We found this argument independently of the SAT computation.
@@ 2.1 Delisle’s symmetry breaking
    Permutation of rows of $A$ and $B$ preserving row equivalence classes (the same permutation applied to $A$ and $B$ simultaneously).
@@
    Permutation of columns of $A$ and $B$ preserving column equivalence classes (the same permutation applied to $A$ and $B$ simultaneously).
@@
    Permutation of the symbols of $A$ preserving $A$-symbol equivalence classes.
@@
    Permutation of the symbols of $B$ preserving $B$-symbol equivalence classes.
@@
Suppose $(A,B)$ is a minimal pair with the symbols in the first column of $A$ not in sorted order within the rows of each row equivalence class. Using row permutations, permute the rows of $(A,B)$ to form $(A^{\prime},B^{\prime})$ such that the rows of the first column of $A^{\prime}$ are now sorted within the rows of each row equivalence class. This is possible because we can freely permute rows of $A$ and $B$ if we focus only on the rows in one particular row equivalence class. Thus, there is some permutation which sorts those particular rows by the entries of the first column of $A$. Consider the first entry $A^{\prime}_{i,0}$ of $A^{\prime}$ that has changed after applying these permutations. (This will also be the first entry of $B^{\prime}$ that has changed, since the same permutations are applied to $A$ and $B$.) Because $A^{\prime}_{i,0}<A_{i,0}$ and $(A_{j,0}^{\prime},B^{\prime}_{j,0})=(A_{j,0},B_{j,0})$ for all $j<i$ we have $(A^{\prime},B^{\prime})<(A,B)$ in contradiction to $(A,B)$ being minimal. ∎
@@
By Proposition 7 if $(A,B)$ is a minimal pair then $A_{0,0}$ must be in $\{0,1,4,7\}$ in case 1 and in $\{0,2,4,6\}$ in cases 3–5. Similarly, by Proposition 8 if $(A,B)$ is a minimal pair then $B_{0,0}$ must be in $\{0,1,4,7\}$ in case 1 and in $\{0,2,4,6\}$ in cases 3–5. Delisle [11, pg. 18] gives the possibilities for the values of $(A_{0,0},B_{0,0})$ in case 1, and the only ones which are in $\{0,1,4,7\}\times\{0,1,4,7\}$ are $(0,0)$, $(1,1)$, $(4,4)$, and $(7,7)$. Similarly, [11, pg. 20] gives the possibilities for the values of $(A_{0,0},B_{0,0})$ in case 3 (and these are identical in cases 4 and 5), and the only ones which are in $\{0,2,4,6\}\times\{0,2,4,6\}$ are $(0,0)$, $(2,2)$, $(4,4)$, and $(6,6)$. ∎
@@ 3.1 Latin square encoding
$x_{i}\mathbin{\rightarrow}s_{i}\quad\text{and}\quad s_{i-1}\mathbin{\rightarrow}s_{i}$ (3)
@@ 3.2 Orthogonality encoding
$(A_{ij}=k\land A_{i^{\prime}j^{\prime}}=k\land B_{ij}=l\land B_{i^{\prime}j^{\prime}}=l)\mathbin{\rightarrow}(i=i^{\prime})$ (4)
for $0\leq i,j,i^{\prime},j^{\prime},k,l<n$ (cf. Zhang [23, Lemma 1]). Taking the contrapositive and writing this using the variables $A_{ijk}$ and $B_{ijk}$, this is equivalent to the clauses $\lnot A_{ijk}\lor\lnot A_{i^{\prime}j^{\prime}k}\lor\lnot B_{ijl}\lor\lnot B_{i^{\prime}j^{\prime}l}$ where $0\leq i,j,i^{\prime},j^{\prime},k,l<n$ with $i\neq i^{\prime}$. This encoding of orthogonality uses $O(n^{6})$ clauses of length 4. An alternative encoding of orthogonality that performs better in practice [23, Lemma 2] uses $O(n^{4})$ clauses of length 3 and $n^{3}$ new auxiliary variables. To describe this orthogonality encoding, we follow the derivation of Bright, Keita, and Stevens [8] based on a composition square.
@@
$(A_{ijk}\land B_{ijl})\mathbin{\rightarrow}Z_{ilk}$ (5)
for $0\leq i,j,k,l<n$. In fact, from $A=ZB$ we also derive the similar clause $(Z_{ilk}\land B_{ijl})\mathbin{\rightarrow}A_{ijk}$, and from $B=Z^{-1}A$ we derive $(Z_{ilk}\land A_{ijk})\mathbin{\rightarrow}B_{ijl}$. These last two types of clauses are logically redundant, but in practice they improve the performance of the SAT solver as they allow the solver to make additional useful propagations.
@@ 3.3 Relation encoding
Similarly, Delisle defines a labelling function $\operatorname{ST\_CLASS}$ on symbol pairs $(s,t)$ where $s$ is a symbol of the first Latin square $A$ and $t$ is a symbol of the second Latin square $B$. Representing $s$ and $t$ as integers in $\{0,\dotsc,9\}$, note that line $s+20$ of the net corresponds to symbol $s$ in the first Latin square, and line $t+30$ of the net corresponds to symbol $t$ in the second Latin square. Concretely, $\operatorname{ST\_CLASS}(s,t)$ is defined to be
@@
$\operatorname{RC\_CLASS}(i,j)=\operatorname{ST\_CLASS}(A_{ij},B_{ij})\text{ for all $0\leq i,j\leq 9$}.$ (8)
That is, if $(A_{ij},B_{ij})=(s,t)$ then $\operatorname{RC\_CLASS}(i,j)=\operatorname{ST\_CLASS}(s,t)$. We encode this condition directly into our SAT encoding. Each case 1–5 is encoded separately, so the forms of $R_{1}$ and $R_{2}$ are known, meaning that the values of $\operatorname{RC\_CLASS}(i,j)$ and $\operatorname{ST\_CLASS}(s,t)$ can be determined in advance for all $0\leq i,j,s,t\leq 9$. We encode the relation constraint in contrapositive form: for all $i$, $j$, $s$, $t$ with $\operatorname{ST\_CLASS}(s,t)\neq\operatorname{RC\_CLASS}(i,j)$ we enforce that $(A_{ij},B_{ij})\neq(s,t)$, i.e., $A_{ij}\neq s$ or $B_{ij}\neq t$. In Boolean logic, this becomes the clauses
$\lnot A_{ijs}\lor\lnot B_{ijt}\text{ where $\operatorname{RC\_CLASS}(i,j)\neq\operatorname{ST\_CLASS}(s,t)$}$ (9)
for all $0\leq i,j,s,t\leq 9$.
@@ 3.4 Symmetry breaking
$\bigwedge_{\begin{subarray}{c}0\leq k,l\leq 9\\ k>l\end{subarray}}(A_{1,0,k}\mathbin{\rightarrow}\lnot A_{0,1,l}).$ (10)
@@
$\bigwedge_{\begin{subarray}{c}0\leq k,l,m\leq 9\\ k\geq l\end{subarray}}\bigl((A_{1,0,m}\land A_{0,1,m}\land B_{1,0,k})\mathbin{\rightarrow}\lnot B_{0,1,l}\bigr).$ (11)
@@
$\bigwedge_{\begin{subarray}{c}i\in R^{\prime}\\ 0\leq l<k\leq 9\end{subarray}}(A_{i,0,k}\mathbin{\rightarrow}\lnot A_{i+1,0,l}).$ (12)
@@
$\bigwedge_{\begin{subarray}{c}j\in C^{\prime}\\ 0\leq l<k\leq 9\end{subarray}}(A_{0,j,k}\mathbin{\rightarrow}\lnot A_{0,j+1,l}).$ (13)
@@
$\bigwedge_{\begin{subarray}{c}s\in S^{\prime}\\ 0\leq i^{\prime}<i\leq 9\end{subarray}}(A_{i,0,s}\mathbin{\rightarrow}\lnot A_{i^{\prime},0,s+1}).$ (14)
@@
$\bigwedge_{\begin{subarray}{c}t\in T^{\prime}\\ 0\leq i^{\prime}<i\leq 9\end{subarray}}(B_{i,0,t}\mathbin{\rightarrow}\lnot B_{i^{\prime},0,t+1}).$ (15)
@@
$\bigwedge_{\begin{subarray}{c}0\leq k,l\leq 9\\ k>l\end{subarray}}(A_{1,0,k}\mathbin{\rightarrow}\lnot B_{1,0,l}).$ (16)
@@
$\bigwedge_{\begin{subarray}{c}0\leq k,l,m\leq 9\\ k>l\end{subarray}}\bigl((A_{1,0,m}\land B_{1,0,m}\land A_{2,0,k})\mathbin{\rightarrow}\lnot B_{2,0,l}\bigr).$ (17)
@@ 4.1 Exhaustive enumeration
$\lnot\Bigl(\bigwedge_{0\leq i,j\leq 9}(A_{i,j,S_{i,j}}\land B_{i,j,T_{i,j}})\Bigr),$ (18)
@@ 5 Results
We now discuss our computational results enumerating all 4-nets(10) with at least two nontrivial relations. Our results were run on an Intel i7 CPU running at 2.8 GHz and using the SAT solver CaDiCaL 1.9.4 using up to 250 MiB of memory. Our scripts are freely available and archived on Zenodo at doi.org/10.5281/zenodo.17352786.
We use a Python script to generate SAT instances in each of the five possible forms (cases 1–5) following the encoding described in Section 3. In order to mitigate the effect of randomness in the search, each case was independently solved 45 times, each time with a different random seed. The differing random seeds ensure that each instance of CaDiCaL makes different choices during the solving process. A tabular summary of the results of these trials is provided in Table 1, and a box plot of the running times is provided in Figure 1.
The SAT solver determined that cases 2, 3, and 4 all had no solutions and these cases were always solvable in a few seconds. It is interesting to note that Delisle [11] ruled out cases 2 and 4 theoretically using a counting argument, but despite the fact we did not explicitly use this in the SAT encoding, the SAT solver quickly proved unsatisfiability on its own. The fact that the solver also quickly ruled out case 3 suggested to us that case 3 was also resolvable using a counting argument, and we were successful in finding one (included at the end of this section). Unfortunately, the proofs produced by the SAT solver, while logically correct, are not intended to be human-readable and the contents of the proofs did not provide us with any mathematical insight.
@@
(a) Cases 1 and 5
(b) Cases 2, 3, and 4
Figure 1: A box plot visualization of the running times of the 45 instances solved in each case.
@@
The fact that the SAT solver was able to quickly rule out case 3 inspired us to look for a counting argument that could rule out this case. We were successful using an approach similar to the arguments used by Gill and Wanless to count the number of points of certain type in a net [13, Thm. 3.1].
@@
The counting argument in the proof of Theorem 5.1 provides some theoretical conditions speculated on by Delisle in their original work: